OpenStudy (anonymous):

find volume of sphere w/ SA of 144 sq. in.

OpenStudy (shane_b):

The formulas you need are:\[\large SA_{sphere}=4\pi r ^2\]\[\large V_{sphere}=\frac{4}{3}\pi r^3\] So you have: \[\large 144in^2=4\pi r ^2\] Solving for r you get: \[\large r=\sqrt{\frac{144in^2}{4 \pi}}=3.385in\]Now just calculate the volume using the radius you just found.